# Wasserstein Stability of Persistence Landscapes and Barcodes

### v1

@@ Abstract
Barcodes form a complete set of invariants for interval decomposable persistence modules and are an important summary in topological data analysis. The set of barcodes is equipped with a canonical one-parameter family of metrics, the $p$-Wasserstein distances. However, the $p$-Wasserstein distances depend on a choice of a metric on the set of interval modules, and there is no canonical choice. One convention is to use the length of the symmetric difference between 2 intervals, which equals to the 1-norm of the difference between their Hilbert functions. We propose a new metric for interval modules based on the rank invariant instead of the dimension invariant. Our metric is topologically equivalent to the metrics induced by the $p$-norms on $\mathbb{R}^{2}$. We establish stability results from filtered CW complexes to barcodes, as well as from barcodes to persistence landscapes. In particular, we show that vectorization via persistence landscapes is 1-Lipschitz with a sharp bound, with respect to the 1-Wasserstein distance on barcodes and the 1-norm on persistence landscapes.
@@ 1 Introduction
The $p$-Wasserstein distances between persistence diagrams (PDs) depend on a choice of metric on the points in the PDs, which represent the interval modules in the interval decomposition. In Section 3, we introduce a new metric on the set of interval modules, denoted by $d_{\operatorname{rank}}$, which compares modules via their rank invariants. Standard choices of metrics for interval modules depend on the dimension invariant [6, 12, 33]. Our metric puts more weight on more persistent features by imposing a higher cost to move points further from the diagonal. The standard choices of metrics induce the same topology on $\mathbb{R}^{2}$ and are uniformly equivalent, whereas our metric $d_{\operatorname{rank}}$ is topologically equivalent while not uniformly equivalent to them. We will use the 1-Wasserstein distance together with $d_{\operatorname{rank}}$, which we denote $W_{1}^{\operatorname{rank}}$. Observe that for any metric $d$ and $p\leq q$, $W_{p}^{d}\geq W_{q}^{d}$, i.e. the 1-Wasserstein distance is the most discriminative distance among all $p$. We equip the set of PDs with $W_{1}^{\operatorname{rank}}$ and show stability from filtered chain complexes to PDs, as well as from PDs to persistence landscapes (PLs).
@@
In practice, one often needs to perform further analysis on topological summaries like PDs, using machine learning and statistical tools which take vectors as inputs. The space of PDs admits metric space structures but not vector space structure. Thus mapping into vector summaries is useful for such analyses. In Section 5, we consider the vectorization using persistence landscapes defined in [3] and show a sharp 1-Lipschitz stability result using the $W_{1}^{\operatorname{rank}}$ on PDs and $L^{1}$ norm on PLs.
In Section 6, we consider an alternative construction of PLs that applies the Möbius inversion of graded rank functions. This produces a sequence of signed PDs, called graded PDs [2]. We use the notion of graded Wasserstein distances to provide a stability result of persistence landscapes in Proposition 6.4.
@@ 1.1 Wasserstein distances and metric for interval modules
    The bottleneck distance is defined using $d_{\infty}$. Some also use $d_{\infty}$ for $p$-Wasserstein distances [1, 30, 20].
@@
    $d_{\operatorname{dim}}(I,J)=\text{length}(I\Delta J)=\int_{\mathbb{R}}\operatorname{dim}I-\operatorname{dim}J|$ [6]
@@ 1.3.1 Related work on barcode stability
There are a number of stability results for PDs based on different constructions. Cohen-Steiner et al. [13] studied PDs arising from Lipschitz functions on compact metric spaces and proved bottleneck stability. Gulen, McCleary, and Patel [17, 23] investigated stability with respect to the edit distance; in particular, Gulen and Patel [17] considered weighted PDs obstained from metric spaces equipped with fully supported probability measures. Both works employed functorial constructions of PDs and used functoriality to establish stability. Other results on stability from filtration functions to PDs include the work of Skraba et al. [33], which generalizes the bottleneck stability to p-Wasserstein stability by choosing $\ell^{p}$ as the ground metric for $W_{p}$. In section 4.1, we rephrase this stability result expressed explicitly by dimension invariants.
Let $f:\sigma\mapsto d_{\operatorname{dim}}(I_{w(\sigma)},I_{v(\sigma)})$ where $I_{w(\sigma)},I_{v(\sigma)}$ are free modules generated by $\sigma$ with respect to $w,v$. Observe f=wvf=w-v. We can rewrite the bottleneck stability[12] as:
@@
The $p$-Wasserstein stability of barcodes in [33] can be reformulated as:
@@ 1.3.2 Related work on vectorization stability
Bubenik proved stability results for PLs [3], using the bottleneck distance on PDs and $L^{\infty}$ norm on PLs. We provide a sharp 1-Lipschitz stability result using $W_{1}^{\operatorname{rank}}$ and the $L^{1}$ norm instead. Various vectorizations of barcodes have been proposed and proven stable [28, 32, 1, 7, 20]. Persistence images, Persistence scale space kernel, tropical coordinates [1, 30, 20] are shown to be Lipschitz w.r.t. $W_{1}$, all with ground metric $d_{\infty}$, and the Lipschitz constants depend on external parameters such as kernels, weights, scales.
@@ 2.2 Persistent homology
Let $I\in\operatorname{Int}(\mathbb{R})$. The corresponding interval module $I$ is an object in $\mathsf{Vect}_{\mathbb{k}}^{\mathbb{R}}$, where $I_{r}=\mathbb{k}$ for $r$ in the interval $I$ and 0 otherwise, and $I_{r\leq s}=1_{\mathbb{k}}$ if $r\leq s$ and $r,s\in I$, and 0 otherwise. We say a persistence module $M$ is interval decomposable if $M\cong\underset{j}{\oplus}I_{j}$ where each $I_{j}$ is an interval module. Interval modules are indecomposable, meaning there is no isomorphism from an interval module $I$ to $\underset{j}{\oplus}N_{j}$ where $N_{j}\in\mathsf{Vect}^{\mathbb{R}}$. Interval decompositions are unique up to re-indexing.
@@
(Structure Theorem in TDA [14]): Let $M$ be a persistence module. If $M$ is pointwise finite dimensional (p.f.d.), then $M$ is interval decomposable, and the interval decomposition is unique up to isomorphism.
@@ 2.3 Wasserstein distances for graded persistence diagrams
In the appendix, we recall the definition of Möbius inversion and the construction of graded PDs from [2]. This section contains the background knowledge for an alternative stability result using graded PDs in section 6.
@@ 3.1 New metric on interval modules
Figure 3: We plot the balls of radius $r$ centered at $x$ for each of $d_{\operatorname{rank}}$ (red), $d_{\operatorname{dim}}$ (blue), and $d_{\infty}$ (green). The radius increases from left to right.
@@
Figure 5: The ball is defined by 4 curves, corresponding to the 4 cases. In particular when an interval doesn’t contain or being contained by $I_{x}$, the boundary curves are hyperbolas.
@@ 3.2 Topological equivalence of metrics
At $(x_{1}+t,x_{2}),\frac{dy_{2}}{dy_{1}}=-1$. This can also be concluded by symmetry. As illustrated in 6,
@@ 4 Wasserstein stability from filtered chain complexes to barcodes
Since $w,v$ are compatible, there exists a total ordering on $K$ respecting $w,v$. And the points in $\operatorname{Dgm}_{i}(w)$, $\operatorname{Dgm}_{i}(v)$ have the same set of birth and death simplex pairs, denoted by $S_{i}=\{(\sigma,\tau):\sigma,\tau\text{ form a birth-death pair}\}$. Consider a matching $\gamma$ that matches the points in Dgm($w$), Dgm($v$) according to the critical simplices.
@@
Figure 8: On the left we visualize the rank functions of the free modules generated by the birth simplex $\sigma$ w.r.t. $w,v$. On th right we compare $d_{\operatorname{rank}}$ to the $L^{1}$ differnce between rank functions of the free modules generated by the simplices.
@@ 4.1 Comparison with previous stability results
Let $f:\sigma\mapsto d_{\operatorname{dim}}(I_{w(\sigma)},I_{v(\sigma)})$. Observe f=wvf=w-v. We can rewrite the bottleneck stability ([12]) as:
@@
The $p$-Wasserstein stability of barcodes shown in [33] can be reformulated as:
@@ 5.1 $L^{1}$-stability for persistence landscapes
Bubenik has shown vectorization using PLs is injective, invertible, and $L^{\infty}$-stable with respect to the bottleneck distance([3]). We’ll show the vectorization is also stable with respect to the 1-Wasserstein distance with the metric $d_{\operatorname{rank}}$.
@@
$\begin{split}\|\Lambda\alpha-\Lambda\beta\|_{1}&=\int_{\mathbb{N}\times\mathbb{R}}|\lambda(k,t)-\lambda^{\prime}(k,t)|d(k,t)\\ &=\sum_{k=1}^{N}\int_{\mathbb{R}}|\lambda(k,t)-\lambda^{\prime}(k,t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|\lambda(k,t)-\lambda^{\prime}(k,t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|u_{(k)}(t)-v_{(k)}(t)|dt\\ &\overset{(1)}{\leq}\int_{\mathbb{R}}\sum_{k=1}^{N}|u_{k}(t)-v_{k}(t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|\Lambda\{(x_{k})\}(t)-\Lambda\{(y_{k})\}(t)|dt\\ &=\sum_{k=1}^{N}\int_{\mathbb{R}}|\Lambda\{(x_{k})\}(t)-\Lambda\{(y_{k})\}(t)|dt\\ &=\sum_{k=1}^{N}d_{\operatorname{rank}}(x_{k},y_{k})\\ &=W_{1}^{\operatorname{rank}}(alpha,\beta)\end{split}$
@@ 5.2 Landscape stability for countable persistence diagrams
Let $\alpha=\underset{i\geq 1}{x_{i}},\beta=\underset{i\geq 1}{y_{i}}$, and for all $k\in\mathbb{N}$, $\alpha_{k}=\underset{1\leq i\leq k}{x_{i}},\beta_{k}=\underset{1\leq i\leq k}{y_{i}}$.
@@
By 5.4, we can fix NN such that $W_{1}^{\operatorname{rank}}(\alpha,\alpha_{N})<\epsilon,W_{1}^{\operatorname{rank}}(\beta,\beta_{N})<\epsilon$.
@@ 6 Graded 1-Wasserstein distances
We recall the construction of graded PDs from [2] in appendix. In short, PDs can be obtained by taking the Möbius inversion of the rank functions, and graded PDs are obtained by taking the Möbius inversion of the graded rank functions. And the grading is on the value of the rank function, with integral value.
@@ A.2 Construction of graded PDs
Let $\mu$ be the Möbius function on the partially ordered set $(\operatorname{Int}(P),\subset)$, given in example A.3. Viewing PDs as integral functions on $\mathbb{R}_{\leq}^{2}$, the author in [29] expressed PDs as:
@@
where $(r,s)\in\mathbb{R}_{\leq}^{2}$. The authors in [2] defined graded PDs by taking the Möbius inversion of the graded rank function instead.

### v2

@@ Abstract
Barcodes form a complete set of invariants for interval decomposable persistence modules and are an important summary in topological data analysis. The set of barcodes is equipped with a canonical one-parameter family of metrics, the $p$-Wasserstein distances. However, the $p$-Wasserstein distances depend on a choice of a metric on the set of interval modules, and there is no canonical choice. One convention is to use the length of the symmetric difference between 2 intervals, which equals to the 1-norm of the difference between their Hilbert functions. We propose a new metric for interval modules based on the rank invariant instead of the dimension invariant. Our metric is topologically equivalent to the metrics induced by the $p$-norms on $\mathbb{R}^{2}$. This metric is sensitive not only to where features appear, but also to how long they persist. We establish stability results from filtered CW complexes to barcodes, as well as from barcodes to persistence landscapes. In particular, we show that vectorization via persistence landscapes is 1-Lipschitz with a sharp bound, with respect to the 1-Wasserstein distance on barcodes and the 1-norm on persistence landscapes.
@@ 1 Introduction
The $p$-Wasserstein distances between persistence diagrams (PDs) depend on a choice of metric on the points in the PDs, which represent the interval modules in the interval decomposition. In Section 3, we introduce a new metric on the set of interval modules, denoted by $d_{\operatorname{rank}}$, which compares modules via their rank invariants. Standard choices of metrics for interval modules depend on the dimension invariant [barcode, bottleneck, stability]. Our new metric is sensitive not only to where features appear, but also to how long they persist. Our metric puts more weight on more persistent features by imposing a higher cost to move points further from the diagonal. The standard choices of metrics induce the same topology on $\mathbb{R}^{2}$ and are uniformly equivalent, whereas our metric $d_{\operatorname{rank}}$ is topologically equivalent while not uniformly equivalent to them. We will use the 1-Wasserstein distance together with $d_{\operatorname{rank}}$, which we denote $W_{1}^{\operatorname{rank}}$. Observe that for any metric $d$ and $p\leq q$, $W_{p}^{d}\geq W_{q}^{d}$, i.e. the 1-Wasserstein distance is the most discriminative distance among all $p$. We equip the set of PDs with $W_{1}^{\operatorname{rank}}$ and show stability from filtered chain complexes to PDs, as well as from PDs to persistence landscapes (PLs).
@@
In practice, one often needs to perform further analysis on topological summaries like PDs, using machine learning and statistical tools which take vectors as inputs. The space of PDs admits metric space structures but not vector space structure. Thus mapping into vector summaries is useful for such analyses. In Section 5, we consider the vectorization using persistence landscapes defined in [pl] and show a sharp 1-Lipschitz stability result using the $W_{1}^{\operatorname{rank}}$ on PDs and $L^{1}$ norm on PLs.
In Section 6, we consider an alternative construction of PLs that applies the Möbius inversion of graded rank functions. This produces a sequence of signed PDs, called graded PDs [Betthauser:2022aa]. We use the notion of graded Wasserstein distances to provide a stability result of persistence landscapes in Proposition 6.4.
@@ 1.1 Wasserstein distances and metric for interval modules
    The bottleneck distance is defined using $d_{\infty}$. Some also use $d_{\infty}$ for $p$-Wasserstein distances [PI, RHB, Kalisnik:2019aa].
@@
    $d_{\operatorname{dim}}(I,J)=\text{length}(I\Delta J)=\int_{\mathbb{R}}\operatorname{dim}I-\operatorname{dim}J|$ [barcode]
@@ 1.3.1 Related work on barcode stability
There are a number of stability results for PDs based on different constructions. Cohen-Steiner et al. [Cohen-Steiner:2010aa] studied PDs arising from Lipschitz functions on compact metric spaces and proved bottleneck stability. Gulen, McCleary, and Patel [gulen2025, MP2022] investigated stability with respect to the edit distance; in particular, Gulen and Patel [gulen2025] considered weighted PDs obstained from metric spaces equipped with fully supported probability measures. Both works employed functorial constructions of PDs and used functoriality to establish stability. Other results on stability from filtration functions to PDs include the work of Skraba et al. [stability], which generalizes the bottleneck stability to p-Wasserstein stability by choosing $\ell^{p}$ as the ground metric for $W_{p}$. In section 4.1, we rephrase this stability result expressed explicitly by dimension invariants.
Let $f:\sigma\mapsto d_{\operatorname{dim}}(I_{w(\sigma)},I_{v(\sigma)})$ where $I_{w(\sigma)},I_{v(\sigma)}$ are free modules generated by $\sigma$ with respect to $w,v$. Observe f=wvf=w-v. We can rewrite the bottleneck stability[bottleneck] as:
@@
The $p$-Wasserstein stability of barcodes in [stability] can be reformulated as:
@@ 1.3.2 Related work on vectorization stability
Bubenik proved stability results for PLs [pl], using the bottleneck distance on PDs and $L^{\infty}$ norm on PLs. We provide a sharp 1-Lipschitz stability result using $W_{1}^{\operatorname{rank}}$ and the $L^{1}$ norm instead. Various vectorizations of barcodes have been proposed and proven stable [PHCetal, RKP, PI, COO, Kalisnik:2019aa]. Persistence images, Persistence scale space kernel, tropical coordinates [PI, RHB, Kalisnik:2019aa] are shown to be Lipschitz w.r.t. $W_{1}$, all with ground metric $d_{\infty}$, and the Lipschitz constants depend on external parameters such as kernels, weights, scales.
@@ 2.2 Persistent homology
Let $I\in\operatorname{Int}(\mathbb{R})$. The corresponding interval module $I$ is an object in $\mathsf{Vect}_{\mathbb{k}}^{\mathbb{R}}$, where $I_{r}=\mathbb{k}$ for $r$ in the interval $I$ and 0 otherwise, and $I_{r\leq s}=1_{\mathbb{k}}$ if $r\leq s$ and $r,s\in I$, and 0 otherwise. We say a persistence module $M$ is interval decomposable if $M\cong\underset{j}{\oplus}I_{j}$ where each $I_{j}$ is an interval module. Interval modules are indecomposable, meaning there is no isomorphism from an interval module $I$ to $\overset{N}{\underset{j}{\bigoplus}}M_{j}$ where $M_{j}\in\mathsf{Vect}^{\mathbb{R}}$ is non-trivial and $N>0$ . Interval decompositions are unique up to re-indexing.
@@
(Structure Theorem in TDA [crawley]): Let $M$ be a persistence module. If $M$ is pointwise finite dimensional (p.f.d.), then $M$ is interval decomposable, and the interval decomposition is unique up to isomorphism.
@@ 2.3 Wasserstein distances for graded persistence diagrams
In the appendix, we recall the definition of Möbius inversion and the construction of graded PDs from [Betthauser:2022aa]. This section contains the background knowledge for an alternative stability result using graded PDs in section 6.
@@ 3.1 New metric on interval modules
Equations 3.1, 3.2 define lines of constant persistence. Equations 3.3, 3.4 define hyperbolas. For different radius, the convexity of the two hyperbolas can change. In the edge case $r=d_{\operatorname{rank}}(x,0)$ where $x$ is the center of the ball, each of the hyperbolas is degenerate, intersecting at a point on the diagonal.
Figure 4: The metric ball’s unconventional shape is due to $d_{\operatorname{rank}}$ uses not the dimension invariants but the rank invariants, dependent on the persistence of the center. The boundaries are always given by 2 hyperbolas and 2 lines of constant persistence. In the edge case $r=d_{\operatorname{rank}}(x,0)$, each of the hyperbolas is degenerate, intersecting at a point on the diagonal.
@@ 3.2 Topological equivalence of metrics
At $(x_{1}+t,x_{2}),\frac{dy_{2}}{dy_{1}}=-1$. This can also be concluded by symmetry. As illustrated in 5,
@@ 4 Wasserstein stability from filtered chain complexes to barcodes
Since $w,v$ are compatible, there exists a total ordering on $K$ respecting $w,v$. And the points in $\operatorname{Dgm}_{i}(w)$, $\operatorname{Dgm}_{i}(v)$ have the same set of birth and death simplex pairs, denoted by $S_{i}=\{(\sigma,\tau):\sigma,\tau\text{ form a birth-death pair}\}$.
Consider a matching $\gamma$ that matches the points in Dgm($w$), Dgm($v$) according to the critical simplices. Let $m$ denote the Lebesgue measure on $\mathbb{R}^{2}$.
@@
Figure 7: On the left we visualize the rank functions of the free modules generated by the simplex $\sigma$ w.r.t. weights $w,v$. On the right we compare the $d_{\operatorname{rank}}$ distance between diagram points with persistence pair $(\sigma,\tau)$ to the $L^{1}$ differnce between rank functions of the free modules generated by $\sigma,\tau$, w.r.t. compatible weights $w,v$.
@@ 4.1 Comparison with previous stability results
Let $f:\sigma\mapsto d_{\operatorname{dim}}(I_{w(\sigma)},I_{v(\sigma)})$. Observe f=wvf=w-v. We can rewrite the bottleneck stability ([bottleneck]) as:
@@
The $p$-Wasserstein stability of barcodes shown in [stability] can be reformulated as:
@@ 5.1 $L^{1}$-stability for persistence landscapes
Bubenik has shown vectorization using PLs is injective, invertible, and $L^{\infty}$-stable with respect to the bottleneck distance([pl]). We’ll show the vectorization is also stable with respect to the 1-Wasserstein distance with the metric $d_{\operatorname{rank}}$.
@@
$\begin{split}\|\Lambda\alpha-\Lambda\beta\|_{1}&=\int_{\mathbb{N}\times\mathbb{R}}|\lambda(k,t)-\lambda^{\prime}(k,t)|d(k,t)\\ &=\sum_{k=1}^{N}\int_{\mathbb{R}}|\lambda(k,t)-\lambda^{\prime}(k,t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|\lambda(k,t)-\lambda^{\prime}(k,t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|u_{(k)}(t)-v_{(k)}(t)|dt\\ &\overset{(1)}{\leq}\int_{\mathbb{R}}\sum_{k=1}^{N}|u_{k}(t)-v_{k}(t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|\Lambda\{(x_{k})\}(t)-\Lambda\{(y_{k})\}(t)|dt\\ &=\sum_{k=1}^{N}\int_{\mathbb{R}}|\Lambda\{(x_{k})\}(t)-\Lambda\{(y_{k})\}(t)|dt\\ &=\sum_{k=1}^{N}d_{\operatorname{rank}}(x_{k},y_{k})\\ &=W_{1}^{\operatorname{rank}}(\alpha,\beta)\end{split}$
@@ 5.2 Landscape stability for countable persistence diagrams
Let $\alpha=\underset{i\geq 1}{\sum}x_{i},\beta=\underset{i\geq 1}{\sum}y_{i}$, and for all $k\in\mathbb{N}$, Let $\alpha_{k},\beta_{k}$ deonte the $k^{\text{th}}$ partial sum of $\alpha,\beta$.
@@
By 5.4, we can fix $N$ such that $W_{1}^{\operatorname{rank}}(\alpha,\alpha_{N})<\epsilon$ and $W_{1}^{\operatorname{rank}}(\beta,\beta_{N})<\epsilon$.
@@ 6 Graded 1-Wasserstein distances
We recall the construction of graded PDs from [Betthauser:2022aa] in appendix. In short, PDs can be obtained by taking the Möbius inversion of the rank functions, and graded PDs are obtained by taking the Möbius inversion of the graded rank functions. And the grading is on the value of the rank function, with integral value.
@@ A.2 Construction of graded PDs
Let $\mu$ be the Möbius function on the partially ordered set $(\operatorname{Int}(P),\subset)$, given in example A.3. Viewing PDs as integral functions on $\mathbb{R}_{\leq}^{2}$, the author in [Patel:2018aa] expressed PDs as:
@@
where $(r,s)\in\mathbb{R}_{\leq}^{2}$. The authors in [Betthauser:2022aa] defined graded PDs by taking the Möbius inversion of the graded rank function instead.